### v1

@@ Abstract
We propose a novel method for fitting machine-learning interatomic potentials with magnetic degrees of freedom, namely, magnetic Moment Tensor Potentials (mMTP). The main feature of the methodology consists in fitting mMTP to magnetic forces (negative derivatives of energies with respect to magnetic moments) derived from spin-polarized density functional theory calculations. We test our method on the bcc Fe-Al system with different composition. Specifically, we calculate formation energies, equilibrium lattice parameter, and total cell magnetization. Our findings demonstrate a precise match between values calculated with mMTP and those obtained by DFT at zero temperature. Additionally, using molecular dynamics, we estimate the finite temperature lattice parameter and capture the cell expansion as was previously revealed in experiment. We demonstrate that mMTPs fitted to magnetic forces, increase the relaxation reliability, which is the percent of successfully relaxed structures (i.e. with almost zero forces, stresses, and magnetic moments after the optimization of geometry). Eventually, we show that the proposed methodology can provide an accurate and reliable mMTP with reduced number of computationally complex spin-polarized density functional theory calculations.
@@ Introduction
Machine-learning interatomic potentials (MLIPs) have become a reliable and computationally efficient tool for materials modeling. [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. Over the past few years a remarkable progress has been achieved in the development of magnetic MLIPs [12, 13, 14, 15, 16, 17, 18, 19, 20], whose functional form explicitly includes magnetic degrees of freedom. These models can be used to predict magnetic properties of materials such as magnetization, Curie temperature, Neel temperature, magnon dispersion, phonon spectra and elastic constants in different magnetic states.
@@
The training set for fitting magnetic MLIPs is typically obtained from spin-polarized density functional theory (DFT) calculations. However, such calculations bear even higher computational burdening than non spin-polarized DFT due to explicit evaluation of magnetic moments of atoms. Besides, accurate training of magnetic MLIPs require a significantly larger training set due to doubling of magnetic MLIPs parameters ($6N$ against $3N$ in non-magnetic MLIPs). Therefore, reduction of the training set size for an efficient fitting of accurate magnetic MLIPs is an urgent problem.
@@
Traditionally, non-magnetic MLIPs are trained to energies, forces and, optionally, stresses, derived from DFT and reproduce these quantity. However, with spin-polarized DFT calculations, we can obtain an additional quantity, which is a negative derivative of energy with respect to magnetic moments, known as magnetic forces. Therefore, to obtain a robust magnetic MLIPs, they can be also trained to reproduce this quantity. The opportunity was first explored by Yuan et al. [20] where the authors fitted equivariant neural network force field (NNFF) to energies, forces, and magnetic forces and accurately predicted magnon dispersion of CrI3 monolayer with the obtained model. More importantly, an accurate NNFF fitting required a relatively small training set size ($<1000$ configurations). This result demonstrates that fitting magnetic MLIPs to magnetic forces can be a possibility to obtain an accurate model using a limited set of spin-polarized DFT calculations.
@@
In this work, we propose a novel method for fitting a magnetic MLIP, specifically, magnetic Moment Tensor Potential (mMTP) [16, 18]. The main principle consists in fitting mMTPs to magnetic forces in addition to energies, forces, and stresses. The training set is obtained using the recently developed constrained DFT (cDFT) method [21] allowing to impose hard constraints on magnetic moments and calculate magnetic forces. Additionally, we employ non-constrained DFT calculations for obtaining the configurations with the equilibrium magnetic moments (i.e., zero magnetic forces). The resulting training set includes 2632 16-atom bcc Fe-Al configurations with different concentrations of Al atoms (up to 50 %) and different random distributions of the atom types within the supercell. Using the obtained training set, we construct two ensembles of mMTPs using two methods. The first method consists in training mMTPs only on energies, forces, and stresses [16, 18]. In the second method, presented in this work, we additionally fit mMTPs to magnetic forces. Further, the two methods are compared in terms of the predictive ability of the fitted mMTPs. Specifically, we predict formation energies, lattice parameters, and total magnetic moment of the unit cell for different concentrations of Al in the Fe-Al system. Additionally, we employ mMTPs trained on magnetic forces in finite temperature molecular dynamics simulation and reproduce the cell expansion trend. Finally, we assess the reliability of the obtained models by the percent of configurations with successfully equilibrated magnetic moments (i.e. almost zero magnetic forces) and configurations with fully relaxed geometry. We note that mMTPs may have up to several thousand of parameters and thus the training set of 2632 configurations is not very large for fitting mMTPs. We show that the novel method produces systematically more accurate and reliable mMTPs on the same training set in comparison with the previous method.
@@
The paper is organized as follows. In “Methods” section we discuss the concept of magnetic MTP and the method of fitting mMTP to magnetic forces. In “Results and Discussion” we first describe training and validation procedures of the fitted ensembles of potentials. After, we perform a quantitative analysis by comparing formation energies, lattice parameters, and total magnetic moments of the unit cell calculated by two ensembles of mMTPs and DFT. In “Conclusion” section we discuss the results of this study.
@@ 1.1 Magnetic Moment Tensor Potential
We denote all the mMTP parameters by $\bm{\theta}=\{\xi_{\alpha},c_{\mu,z_{i},z_{j}}^{\zeta,\beta,\gamma}\}$ and the total energy of configuration by $E=E(\bm{\theta},\bm{R},M)$, where $\bm{R}=({\bm{r}}_{1},\ldots,{\bm{r}}_{n})$ are atomic positions and $M=(m_{1},\ldots,m_{n})$ are magnetic moments of atoms in a configuration. The magnetic MTP is invariant with respect to inversion of magnetic moments and the total energy of configuration calculated by mMTP is:
@@
From Eq. (6) we conclude that the total energy of configuration calculated with mMTP depends on the magnetic moments of the atoms explicitly. Each step of any atomistic simulation with mMTP, e.g. geometry optimization or molecular dynamics simulations, is conducted in two stages. At the first stage, we equilibrate magnetic moments (i.e. minimize energy with respect to magnetic moments) for the fixed atomic positions and lattice vectors. At the second stage, we move atoms and, optionally, change lattice vectors for the fixed magnetic moments. Thus, the correct and reliable equilibration of magnetic moments is critical for mMTP. However, this requires a correct description of magnetic forces, which is discussed in the next section.
@@ 1.2 Fitting to magnetic forces
where $w_{e},w_{f},w_{s}$, and $w_{t}$ are non-negative weights that determine the relative importance of energies, forces, stresses, and magnetic forces during mMTP fitting. Selection of the optimal weights is discussed in the next section. We also demonstrate that minimizing the objective function (8) with the optimal weight $w_{t}$ gives much more accurate magnetic forces than in the case when $w_{t}=0$ (i.e., we do not fit mMTP to magnetic forces). Finally, we show that the method including fitting to magnetic forces is useful for the case of small training set, like the one we use in this work.
@@ 2.1 Training set
For constructing the training set we start from 21 configurations with fully optimized geometries. Among them are 16-atom bcc Fe-Al supercells with different concentrations of Al atoms (from 0 % to 50 %) and different distributions of the atom types within the supercell. Some examples are represented in Fig. 1. The entire training set of 2632 configurations is obtained by randomly perturbing the initial 21 configurations and using 300 K molecular dynamics simulations. Thus, about 80 % of the new configurations are the ones with random perturbation of atomic positions, lattice vectors, and magnetic moments of the initial configurations as described in [18] and the remaining configurations correspond to 300 K molecular dynamics. Further, this training set is used to test our methodology.
@@
The energies, forces, stresses, and magnetic forces for given atomic positions are calculated with DFT and the constrained DFT (cDFT) method [21] implemented in ABINIT [24, 25]. We use a 6 $\times$ 6 $\times$ 6 k-point mesh with a cut-off energy of 25 Hartree and projector augmented wave (PAW) pseudopotentials, implemented within the PBE (generalized gradient approximation) DFT functional. Importantly, about 85 % of the configurations contain non-equilibrium magnetic moments and therefore non-zero magnetic forces.
@@ 2.2 Optimal weight in the case of fitting to magnetic forces
Once we included a magnetic force optimization in the objective function (Eq. (8)), we are required to select an optimal weight for fitting to magnetic forces. To determine the optimal $w_{t}$ value we fix the rest of the weights $w_{e}=1$, $w_{f}=0.01\leavevmode\nobreak\ \mbox{\AA}^{2}$, and $w_{s}=0.001$. Such choice for energy, force, and stress weights is dictated by the previous works where these values were used to accurately fit both non-magnetic and magnetic MTP [26, 16, 18]. We further train mMTPs with different magnetic force weights in the range from $10^{-4}\leavevmode\nobreak\ \mu_{B}^{2}$ to $10^{1}\leavevmode\nobreak\ \mu_{B}^{2}$ and with $N_{\psi}=2$ (415 parameters). We use 5-fold cross-validation to verify the quality of the fitted potentials. The results are represented in Fig.  2.
Figure 2: Validation root mean square errors (RMSEs) for energies, forces, stresses, and magnetic forces with different magnetic force weights in the objective function. Potentials were trained using a 5-fold cross-validation algorithm. We provide the results with 68% confidence interval (i.e. 1-$\sigma$ interval).
From Fig. 2 we can observe a decrease of root mean square error (RMSE) for energy with the increase of magnetic force weight up to $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$. The errors for forces and stress components rise monotonously with the increase of $w_{t}$. At the same time, magnetic force errors expectedly decrease when $w_{t}$ is higher. Analyzing the obtained results we suggest 0.1 $\mu_{B}^{2}$ to be the optimal weight for magnetic forces, as it provides minimal energy errors while force and stress errors are still acceptable.
@@ 2.3 Fitting errors for different magnetic Moment Tensor Potentials
After obtaining the optimal weight for magnetic forces, we compare the accuracy of the potentials fitted on the Fe-Al training set including 2632 configurations. We take different numbers of magnetic basis functions $N_{\psi}$, namely, 2, 3, 4, and 5. We also compare two fitting strategies: without fitting to magnetic forces, i.e. $w_{t}=0\leavevmode\nobreak\ \mu_{B}^{2}$, and with the optimal magnetic force weight $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$. For all the potentials we choose $N_{\phi}=8$, $R_{\rm cut}=4.5\leavevmode\nobreak\ \mbox{\AA}$ , $R_{\rm min}=2.1\leavevmode\nobreak\ \mbox{\AA}$, $M_{\rm max}^{\rm Fe}=3.7\leavevmode\nobreak\ \mu_{B}$, and $M_{\rm max}^{\rm Al}=0.2\leavevmode\nobreak\ \mu_{B}$. The number of parameters for each potential can be found in Tab. 1.
@@
We train an ensemble of five potentials for each number of parameters using the aforementioned training set and then average the fitting errors. After that, we calculate RMSEs between the mMTP predictions and the DFT data. The RMSEs for energy, force, stress, and magnetic force RMSEs are shown in Tab. 2 and visualized in Fig. 3.
Table 2: Root mean square errors for energies, forces, stresses, and magnetic force predicted by mMTPs on the whole training set. $N_{\psi}$ denotes the magnetic basis size and $w_{t}$ is the magnetic force weight. The results are given with 95% confidence interval (i.e. 2-$\sigma$ interval).
@@
From Tab. 2 we see that with an increase in the size of the magnetic basis up to 4, the RMSEs for energy, force, and stress decrease both for zero and non-zero magnetic force weight. This trend can be explained by the higher number of mMTP parameters when using a larger magnetic basis (see Tab. 1). Potentials with $N_{\psi}=5$ have comparable accuracy and sometimes are even less accurate, than those with $N_{\psi}=4$. Thus, in this work we focus on the magnetic basis sizes of 2, 3 and 4 in all following tests. We note that the RMSE for magnetic force decreases significantly when increasing magnetic basis size only if the weight on magnetic forces is non-zero. Hence, the quality of magnetic forces prediction does not affected by magnetic basis size if mMTPs are not fitted to magnetic forces. As it is expected, with $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$ the RMSE for magnetic force is significantly decreased (by a factor of ten) for each number of magnetic basis functions in comparison with the case of $w_{t}=0\leavevmode\nobreak\ \mu_{B}^{2}$. Moreover, fitting to magnetic forces does not significantly increase energy, force, and stress RMSEs. Thus, the mMTP fitting method including training on magnetic forces is promising in terms of accuracy and the predictive ability of the fitted mMTPs.

### v3

@@ Abstract
We developed a method for fitting machine-learning interatomic potentials with magnetic degrees of freedom, namely, magnetic Moment Tensor Potentials (mMTP). The main feature of our method consists in fitting mMTP to magnetic forces (negative derivatives of energies with respect to magnetic moments) as obtained spin-polarized density functional theory calculations. We test our method on the bcc Fe-Al system with different compositions. Specifically, we calculate formation energies, equilibrium lattice parameter, and total cell magnetization. Our findings demonstrate an accurate correspondence between the values calculated with mMTP and those obtained by DFT at zero temperature. Additionally, using molecular dynamics, we estimate the finite-temperature lattice parameter and capture the cell expansion as was previously revealed in experiment. Furthermore, we demonstrate that fitting to magnetic forces increases the reliability of structure relaxation (or, equilibration), in the sense of ensuring that every relaxation run ends up with a successfully relaxed structure (the failure may otherwise be caused by falsely driving a configuration away from the region covered in the training set).
@@ Introduction
Magnetic materials play a key role in numerous technological frontiers including spintronics, novel medical devices, and sensors. In spintronics, they enable devices like magnetic tunnel junctions and spin transistors, leading to faster, more efficient electronics. In medical devices, they are used in magnetic resonance imaging (MRI) machines, targeted drug delivery via magnetic nanoparticles, and cancer treatment through hyperthermia. Magnetic materials also enhance sensor technologies, including magnetoresistive sensors for automotive and industrial applications, and Hall effect sensors for measuring magnetic fields. Thus, study and manipulating the properties of these materials is crucial for the further development of technology. Magnetic materials can be investigated experimentally, both at macro and micro-scales. However, certain limitations preclude the precise determination of their properties. For example, preparing clean samples is essential, as even low concentrations of impurities can significantly affect the accuracy of magnetic properties prediction [1]. Additionally, some methods require accurate calibration of experimental equipment and precise control of temperature conditions [2]. Finally, both samples and experimental equipment can be expensive.
Computational modeling can be used for theoretical investigation of materials. One of the most popular methods is density functional theory (DFT) and spin-polarized DFT, used specifically for modeling of magnetic materials. However, both DFT and spin-polarized DFT are computationally expensive and not applicable for describing physical effects observed in systems containing thousands of atoms. For that reason, alternative interatomic interaction models are being actively developed. Machine-learning interatomic potentials (MLIPs) have gained significant attention as a reliable and computationally efficient tool for materials modeling [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. These potentials were applied to solving important problems via atomistic simulation, such as crystal structure prediction, simulating lattice dynamics, investigating defects in materials, and studying diffusion processes. It would have been difficult to investigate the aforementioned problems using DFT. The listed potentials were developed for investigating non-magnetic materials and, consequently, were limited to studying only a single magnetic state. However, to accurately describe various magnetic states of magnetic materials and predict properties like the Curie temperature or Neel temperature, it is essential to explicitly incorporate the magnetic degrees of freedom of atoms, namely, magnetic moments, into the functional form of the interatomic interaction model. We refer to the MLIPs with magnetic moments in the functional form as magnetic MLIPs. Over the past few years a remarkable progress has been achieved in developing such type of models [14, 15, 16, 17, 18, 19, 20, 21, 22]. The review on the existing classical magnetic interatomic potentials and magnetic MLIPs is given in [23]. The descriptive capabilities and the limitations of these potentials are discussed in this paper.
@@
The training set for fitting magnetic MLIPs is typically obtained from spin-polarized density DFT calculations. However, such calculations bear even higher computational burden than non spin-polarized DFT due to explicit evaluation of magnetic moments of atoms. Besides, accurate training of magnetic MLIPs require a significantly larger training set due to doubling of magnetic MLIPs parameters ($6N$ against $3N$ in non-magnetic MLIPs). Therefore, reduction of the training set size for an efficient fitting of accurate magnetic MLIPs is a critical problem.
@@
Traditionally, non-magnetic MLIPs are fitted to energies, forces and, optionally, stresses, derived from DFT. However, with spin-polarized DFT calculations, we can obtain an additional quantity—a negative derivative of energy with respect to magnetic moments, known as magnetic forces. Therefore, to obtain a robust magnetic MLIPs, they can also be trained to reproduce this quantity. The opportunity was first explored by Yuan et al. [22] where the authors fitted an equivariant neural network force field (NNFF) to energies, forces, and magnetic forces and accurately predicted magnon dispersion of CrI3 monolayer with the obtained model. More importantly, an accurate NNFF fitting required a relatively small training set size ($<1000$ configurations). This result demonstrates that fitting magnetic MLIPs to magnetic forces can be an opportunity to obtain an accurate model using a limited set of spin-polarized DFT calculations.
@@
In this work, we develop a method for fitting magnetic Moment Tensor Potential (mMTP) [18, 20] to magnetic forces in addition to energies, forces, and stresses. The training set is obtained using the recently developed constrained DFT (cDFT) method [24] allowing to impose hard constraints on magnetic moments and calculate magnetic forces. Additionally, we employ non-constrained DFT calculations for obtaining the configurations with the equilibrium magnetic moments (i.e., zero magnetic forces). The resulting training set includes 2632 16-atom bcc Fe-Al configurations with different concentrations of Al atoms (up to 50 %) and different random distributions of the atom types within the supercell. Using the obtained training set, we construct two ensembles of mMTPs using two methods. The first method consists in training mMTPs only on energies, forces, and stresses [18, 20]. In the second method, presented in this work, we additionally fit mMTPs to magnetic forces. Further, the two methods are compared in terms of the predictive ability of the fitted mMTPs. Specifically, we predict formation energies, lattice parameters, and total magnetic moment of the unit cell for different concentrations of Al in the Fe-Al system. Additionally, we employ mMTPs trained on magnetic forces in finite temperature molecular dynamics simulation and reproduce the cell expansion trend. Finally, we assess the reliability of the obtained models by the percentage of configurations with successfully equilibrated magnetic moments (i.e., almost zero magnetic forces) and configurations with fully relaxed geometry. We note that the mMTPs may have up to several thousand of parameters and thus the training set of 2632 configurations is not very large for fitting mMTPs. We show that the novel method produces systematically more accurate and reliable mMTPs on the same training set in comparison with the previous method.
@@
The paper is organized as follows. In the Methods section we introduce the concept of magnetic MTP and the method of fitting mMTP to magnetic forces. In Results and Discussion we first describe the training and validation procedures of the fitted ensembles of potentials. Then we perform a quantitative analysis by comparing formation energies, lattice parameters, and total magnetic moments of the unit cell calculated by two ensembles of mMTPs and DFT. In Conclusion we summarize the results of this study.
@@ 1.1 Magnetic Moment Tensor Potential
We denote a collection of all the mMTP parameters by $\bm{\theta}=\{\xi_{\alpha},c_{\mu,z_{i},z_{j}}^{\zeta,\beta,\gamma}\}$ and the total energy of configuration by $E=E(\bm{\theta},\bm{R},M)$, where $\bm{R}=({\bm{r}}_{1},\ldots,{\bm{r}}_{n})$ are atomic positions and $M=(m_{1},\ldots,m_{n})$ are magnetic moments of atoms in a configuration. The magnetic MTP is invariant with respect to inversion of magnetic moments and the total energy of configuration calculated by mMTP is:
@@
From Eq. (6) we conclude that the total energy of configuration calculated with mMTP depends on the magnetic moments of the atoms explicitly. Each step of any atomistic simulation with mMTP, e.g., geometry optimization or molecular dynamics simulations, is conducted in two stages. At the first stage, we equilibrate magnetic moments (i.e., minimize energy with respect to magnetic moments) for the fixed atomic positions and lattice vectors. At the second stage, we move atoms and, optionally, change lattice vectors for the fixed magnetic moments. Thus, the correct and reliable equilibration of magnetic moments is critical for mMTP. However, this requires a correct description of magnetic forces, which is discussed in the next section.
@@ 1.2 Fitting to magnetic forces
where $w_{e},w_{f},w_{s}$, and $w_{t}$ are non-negative weights that determine the relative importance of energies, forces, stresses, and magnetic forces during the mMTP fitting. The optimal parameters $\bm{\theta}$ are obtained numerically by iterative minimization of the objective function (8) with Broyden-Fletcher-Goldfarb-Shanno algorithm [27]. Selection of the optimal weights is discussed in the next section. We also demonstrate that minimizing the objective function (8) with the optimal weight $w_{t}$ gives much more accurate magnetic forces than in the case when $w_{t}=0$ (i.e., when we do not fit mMTP to magnetic forces). Finally, we show that fitting to magnetic forces is useful for the case of small training set, like the one we use in this work.
@@ 2.1 Training set
For constructing the training set we start with 21 configurations with fully optimized geometries. They are 16-atom bcc Fe-Al supercells with different concentrations of Al atoms (from 0 % to 50 %) and different distributions of the atom types within the supercell. Some examples are represented in Fig. 1. The entire training set of 2632 configurations is obtained by randomly perturbing the initial 21 configurations and using molecular dynamics simulations at 300 K. Thus, about 80 % of the new configurations are the ones with random perturbation of atomic positions, lattice vectors, and magnetic moments of the initial configurations as described in [20] and the remaining configurations correspond to 300 K molecular dynamics. Further, this training set is used to test our methodology.
@@
The energies, forces, stresses, and magnetic forces for the given atomic positions are calculated with DFT and the cDFT method [24] implemented in ABINIT [28, 29]. We use a 6 $\times$ 6 $\times$ 6 k-point mesh with a cut-off energy of 25 Hartree and projector augmented wave (PAW) pseudopotentials, implemented within the Perdew-Burke-Ernzerhof (generalized gradient approximation) DFT functional. Importantly, about 85 % of the configurations contain non-equilibrium magnetic moments and therefore non-zero magnetic forces.
@@ 2.2 Optimal weight in the case of fitting to magnetic forces
Once we included a magnetic force optimization in the objective function (Eq. (8)), we are required to select an optimal weight for fitting to magnetic forces. To determine the optimal $w_{t}$ value we fix the rest of the weights $w_{e}=1$, $w_{f}=0.01$  Å2, and $w_{s}=0.001$. Such choice for energy, force, and stress weights is motivated by the previous works where these values were used to accurately fit both non-magnetic and magnetic MTP [30, 18, 20]. We further train mMTPs with different magnetic force weights in the range from $10^{-4}\leavevmode\nobreak\ \mu_{B}^{2}$ to $10^{1}\leavevmode\nobreak\ \mu_{B}^{2}$ and with $N_{\psi}=2$ (415 parameters). We use 5-fold cross-validation to verify the quality of the fitted potentials. The results are represented in Fig. 2.
Figure 2: Training and validation root mean square errors (RMSEs) for energies, forces, stresses, and magnetic forces with different magnetic force weights in the objective function. Potentials were trained using a 5-fold cross-validation algorithm. We provide the results with 68% confidence interval (i.e., 1-$\sigma$ interval).
From Fig. 2 we can observe a decrease of root mean square error (RMSE) for energy with the increase of magnetic force weight up to $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$. The errors for forces and stress components rise monotonously with the increase of $w_{t}$. At the same time, magnetic force errors expectedly decrease when $w_{t}$ is higher. Analyzing the obtained results we suggest 0.1 $\mu_{B}^{2}$ to be an optimal weight for magnetic forces, as it provides minimal energy errors while force and stress errors are still acceptable. Additionally, as it is seen in Fig. 2, fitting and validation errors are close to each other, i.e., the training set was constructed correctly. We further fit mMTPs on the entire training set.
@@ 2.3 Fitting errors for different magnetic Moment Tensor Potentials
We train an ensemble of five potentials for each number of parameters using the aforementioned training set and then average the fitting errors. After that, we calculate RMSEs between the mMTP predictions and the DFT data. The RMSEs for energy, force, stress, and magnetic force RMSEs are shown in Table 2 and visualized in Fig. 3.
Table 2: Root mean square errors for energies, forces, stresses, and magnetic force predicted by mMTPs on the whole training set. $N_{\psi}$ denotes the magnetic basis size and $w_{t}$ is the magnetic force weight. The results are given with 95% confidence interval (i.e., 2-$\sigma$ interval).
@@
From Table 2 we see that with an increase in the size of the magnetic basis up to 4, the RMSEs for energy, force, and stress decrease both for zero and non-zero magnetic force weight. This trend can be explained by the higher number of mMTP parameters when using a larger magnetic basis (see Table 1). Potentials with $N_{\psi}=5$ have comparable accuracy and sometimes are even less accurate than those with $N_{\psi}=4$. Thus, in this work we focus on the magnetic basis sizes of 2, 3 and 4 in all following tests. We note that the RMSE for magnetic force decreases significantly with the increase in magnetic basis size only if the weight on magnetic forces is non-zero. Hence, the quality of magnetic forces prediction does not affected by magnetic basis size if mMTPs are not fitted to magnetic forces. As it is expected, with $w_{t}=0.1\leavevmode\nobreak\ \mu_{B}^{2}$ the RMSE for magnetic force is significantly decreased (by a factor of ten) for each number of magnetic basis functions in comparison with the case of $w_{t}=0\leavevmode\nobreak\ \mu_{B}^{2}$. Moreover, fitting to magnetic forces does not significantly increase energy, force, and stress RMSEs. Thus, the mMTP fitting method including training on magnetic forces is promising in terms of accuracy and the predictive ability of the fitted mMTPs.